OpenStudy (anonymous):

The perimeter of a triangle is twice the sum of its length and its width, The perimeter is 22 meters and its length is 2 meters more than twice its width. What is its length

OpenStudy (anonymous):

P=2(l+w) 22=2((2w+2)+w) 22=4w+4+w 22=5w+4 22-4=18 18=5w 18/5=3.6 L=3.6